OpenStudy (anonymous):

A corner store sells two kinds of baked goods: cakes and pies. A cake costs $5 and a pie costs $7. In one day, the store sold 15 baked goods for a total of $91. How many cakes did they sell? A. 8 cakes B. 7 cakes C. 4 cakes D. 5 cakes

OpenStudy (anonymous):

c+p=15 (multiply by 5) =5c+5P=75 5c+7P=91 subtract 1st from 2nd=> 2p=16 p=16/2=8 substitute p into 1st: c+8=15 c=15-8=7 Therefore 8 pies and 7 cakes Answer is b.7 cakes